A race car goes around a level, circular track with a diameter of 1.00 km at a constant speed of 89 km/h. What is the car's centripetal acceleration in m/s2?
Answer #1

Centripetal acceleration is given by:

ac = V^2/R

V = Speed of car = 89 km/hr = (89 km/hr)*(1000 m/1 km)*(1 hr/3600 sec)

V = 89*1000/3600 = 24.722 m/sec

R = Radius of track = diameter/2 = 1 km/2 = 0.5 km = 500 m

Using these values:

ac = 24.722^2/500

ac = 1.22 m/sec^2 = Centripetal acceleration

Please Upvote.